OpenStudy (anonymous):

Write equation for the following transformation of y=x: a vertical stretch by a factor of 6

OpenStudy (anonymous):

A. y=6x B. y=x/6 C.y=x+6 D. y=x-6

OpenStudy (triciaal):

stretch affects the y value when a = 6 means 6y original y = x 6 y on the left therefore do the same to the right the slope is a constant ratio change in y over the corresponding change in x if the y changes to 6 then to maintain the slope of 1 the change in x has to be 1/6(x)